OpenStudy (anonymous):

Eliminate the parameter. x = 3 cos t, y = 3 sin t I got to the part where its (x/3) = cos t. where do I go from now?

OpenStudy (anonymous):

Now, you want to use: \[ \cos^2(t)+\sin^2(t) = 1 \]

OpenStudy (anonymous):

would I use that in place of cos to cancel out the right side and leave t, and the right side would be cos^2(x/3) = t?

OpenStudy (anonymous):

\[ x^2+y^2=(3\cos t)^2+(3\sin t)^2 \]

OpenStudy (anonymous):

im still confused. :( I just don't get how to eliminate the parameters in the questions with trig functions

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

\[ =9(\cos^2t+\sin^2t) \]The \(t\) will go away

OpenStudy (anonymous):

how would the t go away? could you show me the whole process please with an explanation?:)

OpenStudy (anonymous):

Do you know algebra?

ganeshie8 (ganeshie8):

``` x = 3 cos t, y = 3 sin t I got to the part where its (x/3) = cos t. where do I go from now? ``` how did you get (x/3) = cos t ?

OpenStudy (anonymous):

OH so would the answer be (x^2/9) + (y^2/9) = 1???

ganeshie8 (ganeshie8):

yes thats right! but nobody leaves the equation of circle in that form. you may want to kick that 9 to the right side

ganeshie8 (ganeshie8):

multiply 9 through out

OpenStudy (anonymous):

ok thank you so much!!!! you guys all rock:)

OpenStudy (anonymous):

is it possible to give both of you guys medals??

ganeshie8 (ganeshie8):

Notice, you need to know few trig identities to work these... time to revise your trig

ganeshie8 (ganeshie8):

\[\sin^2t + \cos^2t=1\] you would be using this identity a lot, definitely worth memorizing